OpenStudy (vshiroky):

If Marty plants 5 seeds to get 3 flowers, how many seeds will he need to produce 72 flowers?

OpenStudy (anonymous):

\[5seeds \rightarrow3flowers \] \[x \rightarrow72\]

OpenStudy (anonymous):

then x=(5*72)/3

OpenStudy (vshiroky):

How would I set it up as fractions? would it be 3/5=x/72?

OpenStudy (anonymous):

i think it is (5*72)/3

OpenStudy (vshiroky):

Ok I have another one..

OpenStudy (vshiroky):

If paul can swim 2 laps in 5 minutes, how long will it take him to swim 13 laps?

OpenStudy (anonymous):

\[2\rightarrow5\] \[13rightarrowx\]

OpenStudy (vshiroky):

when you do the arrow what does that mean?

OpenStudy (vshiroky):

@esamalaa

OpenStudy (anonymous):

2 lips (at) 5 mins. 13 lips(at) x mins.

OpenStudy (vshiroky):

Would I do 13*5/2?

OpenStudy (anonymous):

yes